OpenStudy (anonymous):

douald has investments totaling 4,000 in two accounts-one a savings account paying 3% interest and the other bond paying 5%. if the annual from the two investment was 180%, how much did he have in invested at each rate?

OpenStudy (anonymous):

douald has investments totaling 4,000 in two accounts-one a savings account paying 3% interest and the other bond paying 5%. if the annual from the two investment was 180, how much did he have in invested at each rate?

OpenStudy (anonymous):

Let funds deposited into account 1 be A and funds deposited into account 2 be B. A + B = 4000, since the total amount invested is $4,000. Since A earns interest at 3% and B earns interest at 5%, we can set up a second equation like this: A*.03 + B*.05 = 180

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

So you have A+B=4000 And .03A + .05B=180 Can you solve one of those for A or B? The first one, maybe?

OpenStudy (anonymous):

no can u show me how

OpenStudy (anonymous):

Let's solve for A by subtracting B from both sides: A+B=4000 - B -------------- A =4000-B

OpenStudy (anonymous):

Now you can substitute that in for A in the second equation: .03(4000-B) + .05B = 180

OpenStudy (anonymous):

i thought i subtract the like terms

OpenStudy (anonymous):

I'm not sure what you mean.

OpenStudy (anonymous):

o while ill do it your way

OpenStudy (anonymous):

ok wat i do now

OpenStudy (anonymous):

Distribute .03: .03*4000 - .03B + .05B = 180

OpenStudy (anonymous):

ok i did it

OpenStudy (anonymous):

Okay, what did you get? Then simplify it.

OpenStudy (anonymous):

is the answer 120.02

OpenStudy (anonymous):

Well, .03*4000 = 120.00. But that is not the answer to the problem. We need to solve for A and B.

OpenStudy (anonymous):

ok is u workin it out to

OpenStudy (anonymous):

I want you to understand how to get there, I'd rather not just give you the answer.

OpenStudy (anonymous):

When you distribute .03: .03*4000 -.03B +.05B what do you get?

OpenStudy (anonymous):

i know hw answer i just said was the wronge answer

OpenStudy (anonymous):

.03*4000 -.03B+.05B = 120 + .02B = 180 Subtract 120 from both sides .02B=60 Divide 60/.02=B=3000 Plug back in to the original equation A+3000=4000 Therefore A = 1000 and B = 3000

OpenStudy (anonymous):

want it be -0.35b=60

OpenStudy (anonymous):

then wat do i do next

OpenStudy (anonymous):

I worked it all out above.

OpenStudy (anonymous):

how much did he have invested at each rate?

OpenStudy (anonymous):

I already answered the question. .03*4000 -.03B+.05B = 120 + .02B = 180 Subtract 120 from both sides .02B=60 Divide 60/.02=B=3000 Plug back in to the original equation A+3000=4000 Therefore A = 1000 and B = 3000